let the fundamental frquency is 45hz and the length of the wire is 87.5cm. If the linear density of the wire is 4*10^-2.find the tension in the string

Grade:12

1 Answers

Arun
5 years ago
Dear Amal
 
lambda = 2 * l = 2 * 0.875 = 1.75 m
speed of transverse wave v = 45 * 1.75 = 78.75 m / sec
 
Tension = v^2 * m = 78.75^2 * 0.04 = 248.06 N